# Generative-Enhanced Heterogeneous Graph Contrastive Learning

### v2

@@ 3. PRELIMINARIES
Heterogeneous Information Network (HIN): A HIN is characterized by a graph structure $G=(V,E,I,J,\phi,\varphi)$, in which the collections of nodes and edges are symbolized by $V$ and $E$ respectively. For every node $v$ and edge $e$, there are associated type mapping functions $\phi:V\rightarrow I$ for nodes and $\varphi:E\rightarrow J$ for edges. Here, $I$ represents the node types and $J$ denotes the edge types, with the total number of types exceeding two, i.e., $|I|+|J|>2$.
@@
Meta-path: In an $\operatorname{HIN}G=(V,E,I,J,\phi,\varphi)$, a meta-path $\rho$ is represented as $I_{1}\stackrel{{\scriptstyle J_{1}}}{{\longrightarrow}}I_{2}\stackrel{{\scriptstyle J_{2}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle J_{l}}}{{\longrightarrow}}I_{l+1}$, describing a composite connection between $I_{1}$ and $I_{l+1}$. Taking Figure 2 as an instance, multiple meta-paths, like “Paper-Author-Paper (PAP)” and “Paper-Subject-Paper (PSP)” can link entities within the HIN. Commonly, different meta-paths can reveal the different inter-dependency information of two papers. The path “PAP” indicates that two papers were written by the same author, while “PSP” illustrates that two papers are of the same subject.
@@
Figure 3. The training framework of our proposed GC-HGNN. (a) The model respectively takes the entire heterogeneous graph and the views based on meta-paths, dividing them into inputs for the network schema View and the meta-path view; (b) the network schema view employs an attention mechanism to focus on one-hop neighbors of each node; (c) the meta-path view utilizes a masked autoencoder to generate views for intra-contrast; (d) the hierarchical contrast uses a proposed dynamic sampling mechanism to generate more hard negative samples to enhance the contrastive discriminator.
4. THE PROPOSED MODEL
@@
In this section, our overall framework and brief module descriptions are shown in Figure 3. We then provide detailed explanations for each module following the framework’s order. Finally, we propose different optimization strategies based on the downstream tasks.
4.1. Network Schema View
@@
This section typically takes the entire heterogeneous information network as input, as shown in Figure 3 (a). First, we consider each node in the network as an anchor, forming a subnetwork that includes all the one-hop neighbor nodes. For example, in Figure 3 (b), the subgraph indicated by the circles represents the subnetwork composed of the target node. Next, we put such subnetworks into a heterogeneous neighbor aggregation network, utilizing attention mechanisms to obtain node embeddings. Finally, we select all the embeddings of the desired target node for downstream tasks. Specific implementation process: We project all nodes into the same node space ($\boldsymbol{P}_{r}$ is projection matrix) because the subnetworks contain different node types. We define $e$ as the current node computing the attention score and $\boldsymbol{h}_{e}^{0}$ as its initial embedding. Next, we learn specific aggregation functions for different types of edges to compute attention scores:
@@
(1) $\alpha_{ew}=\frac{\exp\left(f_{r}\left(\boldsymbol{h}_{e}^{0},\boldsymbol{P}_{r}\boldsymbol{h}_{w}^{0}\right)\right)}{\sum_{j\in\mathcal{N}_{1}(e)}\exp\left(f_{r}\left(\boldsymbol{h}_{e}^{0},\boldsymbol{P}_{r}\boldsymbol{h}_{j}^{0}\right)\right)},\quad\forall w\in\mathcal{N}_{1}(e)\$
@@
where $f_{r}(\cdot,\cdot)$ denotes the deep neural network designed to execute attention mechanisms over one-hop neighbors. Here, $\alpha_{ew}$ represents the influence level for neighbor node $w$, and $\mathcal{N}_{1}(e)$ denotes the set of one-hop neighbor nodes of $e$. Finally, we aggregate the following information from $\mathcal{N}_{1}(e)$:
(2) $\boldsymbol{h}_{e}^{1}=\sigma\left(\sum_{w\in\mathcal{N}_{1}(e)}\alpha_{ew}\boldsymbol{h}_{w}^{0}\right)$
@@
4.2. Meta-path View
In this part, model inputs are multi-views constructed from different meta-paths. To mitigate the imbalanced data augmentation, we enhance multiple views through a generative mask autoencoder. For example, in Figure 3 (a), we connect two nodes of the same type using the specified meta-path ‘PAP’. Thus, such a graph composed of nodes of the same type is referred to as a view based on the meta-path ‘PAP’. Based on preliminaries, we define a set of meta-paths as $\left\{\rho_{1},\ldots,\rho_{j},\ldots,\rho_{M}\right\}$, and the corresponding views are $\left\{G^{\rho_{1}},G^{\rho_{2}},\ldots,G^{\rho_{M}}\right\}$. $M$ is the number of meta-paths we have selected. Here, the model defines a graph $\mathcal{G}=(\mathcal{V},\boldsymbol{A},\boldsymbol{X})$. For each graph, $\mathcal{V}$ is the node set, $N=$ $|\mathcal{V}|$ is the number of nodes, $\boldsymbol{A}\in\{0,1\}^{N\times N}$ is the adjacency matrix, and $\mathbf{X}\in\mathbb{R}^{N\times d}$ is the input node feature matrix.
First, we employ a non-replacement random sampling strategy to select a set of nodes $\mathcal{V}^{[Mask]}$ from the entire set of nodes $\mathcal{V}$.
(3) $\mathcal{V}^{[Mask]}=\{\mathcal{V}_{i}\in\mathcal{V}\mid r_{i}\leq p\},r_{i}\sim\text{Uniform}(0,1)$
where $p$ is the mask ratio we set. The uniform random sampling (Hou et al., 2022) helps the encoder avoid potential bias centers, which refers to the bias arising when all neighbors of a node are either completely masked or completely visible. We mask the entire node features in the set $\mathcal{V}^{[Mask]}$ to obtain $\widetilde{\boldsymbol{X}}$.
@@ 4.2. Meta-path View
(4) $\boldsymbol{H}=f_{E}^{(L)}(\boldsymbol{A},\widetilde{\boldsymbol{X}})$
where $\widetilde{\boldsymbol{X}}$ is feature input with the $[Mask]$, and $\boldsymbol{H}\in\mathbb{R}^{N\times d}$ denotes code after the encoder. Next, to further encourage the decoder to learn compressed representations, we use a new ratio and the same sampling strategy to mask a different set of nodes $\mathcal{V}^{[ReMask]}$. We use a GNN with the same structure as the encoder to act as the decoder. However, the decoder can be GAT (Veličković et al., 2018), GCN (He et al., 2020a), and its layer number and type can be independent of the encoder’s.
(5) $\boldsymbol{H}_{Re}=f_{D}^{(L)}(\boldsymbol{A},\widetilde{\boldsymbol{H}})$
where $\widetilde{\boldsymbol{H}}$ is decoder input with the $[Remask]$, and $\boldsymbol{H}_{Re}$ as the reconstructed output under meta-path view. We perform multiple masking operations between encoders, equivalent to the masked nodes reconstructing their features through their neighbors. The uniform sampling strategy ensures that each reconstructed node receives the same supervised signals. Such an autoencoder helps train robust node embeddings and mitigates the impact of noise from the meta-path view.
@@
(6) $w_{\rho_{j}}=\frac{1}{|\mathcal{V}|}\sum_{\boldsymbol{e}\in\mathcal{V}}\boldsymbol{q}^{T}\cdot\tanh\left(\boldsymbol{W}{\boldsymbol{h}_{Re}^{\rho_{j}}}+b\right)$
@@
(7) $\gamma_{\rho_{j}}=\frac{\exp\left(w_{\rho_{j}}\right)}{\sum_{j=1}^{N}\exp\left(w_{\rho_{j}}\right)}$
@@
where $\gamma_{\rho_{j}}$ denotes the weight of meta-path $\rho_{j}$. According to the attention mechanism (Wang et al., 2019), when the semantic attention score $\gamma_{\rho_{j}}$ is higher, it indicates the meta-path $\rho_{j}$ gets more importance in downstream tasks. We aggregate feature embeddings from all meta-paths to obtain the final node features under the meta-path view.
(8) $\boldsymbol{h}_{e}^{2}=\sum_{j=1}^{N}\gamma_{\rho_{j}}\cdot\boldsymbol{h}_{e}^{\rho_{j}}$
4.3. Hierarchical Contrastive learning
The inputs for this part are the embeddings obtained from network schema and meta-path: $\boldsymbol{h}_{e}^{1}$ and $\boldsymbol{h}_{e}^{2}$. To alleviate sampling bias and information loss in the contrastive discriminator, we propose novel sampling strategies and hierarchical contrast. The model employs two sampling strategies. For intra-contrast, it uses a node-level contrast, whereas for inter-contrast, it adopts an edge-level contrast. Edge-level sampling is illustrated in Figure 4. Next, we get an edge sampling set $\mathcal{E}$.
@@ 4.3. Hierarchical Contrastive learning
Figure 4. Our GC-HGNN proposed a hard sampling strategy for inter-contrast. Location-aware Pos: the green node uses metapath2vec (Dong et al., 2017) to perceive the location within a heterogeneous graph to obtain node embeddings. Semantic-aware Pos: The blue node is connected to the anchor node through all types of meta-paths, indicating the semantic similarity between them.
@@
Intra-contrast. First, we utilize sampling set $\mathcal{E}$ to select the positive sample set $\widetilde{\mathcal{V}}$ for node-level contrast based on the degree of nodes. High-degree nodes have more connections in the graph and are more likely to capture the central structure and important patterns in the graph. The embeddings of the model, based on multi-views, are denoted by $\left\{\boldsymbol{H}_{Re}^{\rho_{1}},\ldots,\boldsymbol{H}_{Re}^{\rho_{j}},\ldots,\boldsymbol{H}_{Re}^{\rho_{M}}\right\}$. Next, we take these embeddings as input to obtain a multi-view loss. Intra-contrast employs scaled cosine error (SCE) (Grill et al., 2020) as the contrastive loss between multi-views. SCE has been proven to focus on hard samples in the data to improve the stability of representation learning.
@@
(9) $\mathcal{L}_{Intra}=\frac{1}{|\widetilde{\mathcal{V}}|}\sum_{v_{i}\in\widetilde{\mathcal{V}}}\left(1-\frac{\boldsymbol{x}_{i}^{T}\boldsymbol{z}_{i}}{\left\|\boldsymbol{x}_{i}\right\|\cdot\left\|\boldsymbol{z}_{i}\right\|}\right)^{\eta},\eta\geq 1$
@@
Inter-contrast. Inter-contrast performs collaborative contrastive learning between two non-aligned views. In set $\mathcal{E}$, we define the target node to be sampled as $i$ and the connected nodes to it as a positive set $\mathbb{P}_{i}$, inspired by (Wang et al., 2021). Otherwise, they form the negative sample set $\mathbb{N}_{i}$. From the perspective of the scale of $\mathcal{E}$ and $\widetilde{\mathcal{V}}$, inter-contrast typically has more hard negative samples, and thus, its supervisory role is more evident. The contrastive loss of the network schema view with respect to the meta-path view is as follows:
(10) $\mathcal{L}_{Inter}^{1}=-\log\frac{\sum_{j\in\mathbb{P}_{i}}\exp\left(\operatorname{sim}\left(h_{i}^{1},h_{j}^{2}\right)/\tau\right)}{\sum_{k\in\left\{\mathbb{P}_{i}\cup\mathbb{N}_{i}\right\}}\exp\left(\operatorname{sim}\left(h_{i}^{1},h_{k}^{2}\right)/\tau\right)}$
where $\operatorname{sim}(i,j)$ is the cosine similarity, and $\tau$ denotes a temperature parameter. Inter-contrast utilizes mutual contrast between the two views to achieve the collaborative process. The contrastive loss of the meta-path view with respect to the network schema view is $\mathcal{L}_{Inter}^{2}$. The overall inter-contrast losses are as follows:
(11) $\mathcal{L}_{{Inter}}=\frac{1}{|\mathcal{E}|}\sum_{i\in\mathcal{E}}\left[\lambda\cdot\mathcal{L}_{Inter}^{1}+(1-\lambda)\cdot\mathcal{L}_{Inter}^{2}\right]$
where $\lambda$ is a balance coefficient to control the weight of each view pair. We employed intra-contrast and inter-contrast for hierarchical contrastive learning and analyzed their operational mechanisms. The intra-contrast assumes that meta-paths are interconnected and aim to maximize their mutual information (Liu et al., 2021). This method integrates the different semantics across meta-paths to form a generalized knowledge representation for downstream tasks. On the other hand, the inter-contrast approach attempts to enhance node representation by contrasting local and global information, preventing nodes from being influenced by noise from higher-order neighbors. The hierarchical contrast and sampling strategies are designed to construct a useful contrastive discriminator to enhance the adversarial interplay between the generative and contrastive models.
4.4. Model Optimization

### v3

@@ II PRELIMINARIES
Heterogeneous Information Network (HIN): A HIN is characterized by a graph structure $\mathcal{G}=(V,E,I,J,\phi,\varphi)$, in which the collections of nodes and edges are symbolized by $V$ and $E$ respectively. For every node $v$ and edge $e$, there are associated type mapping functions $\phi:V\rightarrow I$ for nodes and $\varphi:E\rightarrow J$ for edges. Here, $I$ represents the node types and $J$ denotes the edge types, with the total number of types exceeding two, i.e., $|I|+|J|>2$.
@@
Meta-path: In an HIN $\mathcal{G}=(V,E,I,J,\phi,\varphi)$, a meta-path $\rho$ is depicted as $I_{1}\stackrel{{\scriptstyle J_{1}}}{{\longrightarrow}}I_{2}\stackrel{{\scriptstyle J_{2}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle J_{l}}}{{\longrightarrow}}I_{l+1}$, which describes a composite connection between $I_{1}$ and $I_{l+1}$. Taking Figure 2 as an instance, multiple meta-paths, such as “Paper-Author-Paper (PAP)” and “Paper-Subject-Paper (PSP)” can link entities within the HIN. Typically, different meta-paths can reveal the varying inter-dependency information between two papers. The meta-path “PAP” indicates that two papers were written by the same author, while “PSP” illustrates that two papers belong to the same subject. Before introducing the proposed model, we summarize the symbols used in Table I in this paper.
@@
Figure 3: The training framework of our proposed GC-HGNN. (a) The model respectively takes the entire heterogeneous graph and the views based on meta-paths, dividing them into inputs for the network schema View and the meta-path view; (b) the network schema view employs an attention mechanism to focus on one-hop neighbors of each node; (c) the meta-path view utilizes a masked autoencoder to generate views for intra-contrast; (d) the hierarchical contrast uses a proposed dynamic sampling mechanism to generate more hard negative samples to enhance the contrastive discriminator.
III THE PROPOSED MODEL
@@
In this section, we present overall framework and brief module descriptions, as shown in Figure 3. We then provide detailed explanations for each module following the framework’s order. Finally, we propose different optimization strategies based on the downstream tasks.
III-A Network Schema View
@@
This section takes the entire heterogeneous information network as input, as shown in Figure 3 (a). First, we consider each node in the network as an anchor, forming a subnetwork that encompasses all the one-hop neighbor nodes. For example, in Figure 3 (b), the subgraph indicated by the circles represents the subnetwork formed by the target node. Next, we input such subnetworks into a heterogeneous neighbor aggregation network, utilizing attention mechanisms to obtain node embeddings. Finally, we select the embeddings of the desired target node for downstream tasks. Specific implementation process: We project all nodes into the same node space ($\boldsymbol{P}_{r}$ is projection matrix) due to the various node types within the subnetworks. We define $e$ as the current node computing the attention score, and $\boldsymbol{h}_{e}^{0}$ as its initial embedding. Next, we learn aggregation functions for different types of edges to compute attention scores:
@@
$\alpha_{ew}=\frac{\exp\left(f_{r}\left(\boldsymbol{h}_{e}^{0},\boldsymbol{P}_{r}\boldsymbol{h}_{w}^{0}\right)\right)}{\sum_{j\in\mathcal{N}_{1}(e)}\exp\left(f_{r}\left(\boldsymbol{h}_{e}^{0},\boldsymbol{P}_{r}\boldsymbol{h}_{j}^{0}\right)\right)},\quad\forall w\in\mathcal{N}_{1}(e)\$ (1)
@@
where $f_{r}(\cdot,\cdot)$ denotes the deep neural network designed to implement attention mechanisms over one-hop neighbors. Here, $\alpha_{ew}$ is the influence level for neighbor node $w$, and $\mathcal{N}_{1}(e)$ denotes the set of one-hop neighbor nodes of $e$. Finally, we aggregate the following information from $\mathcal{N}_{1}(e)$:
$\boldsymbol{h}_{e}^{1}=\sigma\left(\sum_{w\in\mathcal{N}_{1}(e)}\alpha_{ew}\boldsymbol{h}_{w}^{0}\right)$ (2)
@@
III-B Meta-path View with Generative Learning
In this section, the model inputs are multi-views constructed from different meta-paths. To address the issue of imbalanced data augmentation, we enhance multiple views using a generative mask autoencoder. For example, in Figure 3 (a), we connect two nodes of the same type via the specified meta-path ‘PAP’. Thus, such a graph composed of nodes of the same type is termed a view based on the meta-path ‘PAP’. Based on the preliminaries, we define a set of meta-paths as $\left\{\rho_{1},\ldots,\rho_{j},\ldots,\rho_{M}\right\}$, and the corresponding views as $\left\{G^{\rho_{1}},G^{\rho_{2}},\ldots,G^{\rho_{M}}\right\}$. $M$ denotes the number of meta-paths we have selected. Here, the model defines a graph $G=(\mathcal{V},\mathcal{A},\boldsymbol{X})$. For each graph, $\mathcal{V}$ is the node set, $N=$ $|\mathcal{V}|$ is the number of nodes, $\mathcal{A}\in\{0,1\}^{N\times N}$ is the adjacency matrix, and $\mathbf{X}\in\mathbb{R}^{N\times d}$ is the input node feature matrix.
First, we employ a sampling strategy without replacement to select a set of nodes $\mathcal{V}^{[Mask]}$ from the entire nodes set $\mathcal{V}$:
$\mathcal{V}^{[Mask]}=\{\mathcal{V}_{i}\in\mathcal{V}\mid r_{i}\leq p\},r_{i}\sim\text{Uniform}(0,1)$ (3)
where $p$ is the mask ratio we set, and $\mathcal{V}_{i}$ is the node set based on $r_{i}$. The uniform random sampling [28] helps the encoder avoid potential bias centers, which refers to the bias arising when all neighbors of a node are either completely masked or completely visible. We mask the entire node features in the set $\mathcal{V}^{[Mask]}$ to obtain $\widetilde{\boldsymbol{X}}$.
@@ III-B Meta-path View with Generative Learning
$\boldsymbol{H}=f_{E}^{(L)}(\mathcal{A},\widetilde{\boldsymbol{X}})$ (4)
where $\widetilde{\boldsymbol{X}}$ is feature input with the $[Mask]$, and $\boldsymbol{H}\in\mathbb{R}^{N\times d}$ denotes code after the encoder. Next, to further encourage the decoder to learn compressed representations, we use a new ratio and the same sampling strategy to mask a different set of nodes $\mathcal{V}^{[ReMask]}$. We use a GNN with the same structure as the encoder to act as the decoder. However, the decoder can be GAT [29], GCN [30], and its layer number and type can be different.
$\boldsymbol{H}_{Re}=f_{D}^{(L)}(\mathcal{A},\widetilde{\boldsymbol{H}})$ (5)
where $\widetilde{\boldsymbol{H}}$ is decoder input with the $[Remask]$, and $\boldsymbol{H}_{Re}$ as the reconstructed output under meta-path view. We perform multiple masking operations between encoders, equivalent to the masked nodes reconstructing their features through their neighbors. The uniform sampling strategy ensures that each reconstructed node receives the same supervised signals. Such an autoencoder helps train robust node embeddings and mitigates the impact of noise.
@@
$w_{\rho_{j}}=\frac{1}{|\mathcal{V}|}\sum_{\boldsymbol{e}\in\mathcal{V}}\boldsymbol{q}^{T}\cdot\tanh\left(\boldsymbol{W}{\boldsymbol{h}_{Re}^{\rho_{j}}}+b\right),$ (6)
@@
$\gamma_{\rho_{j}}=\frac{\exp\left(w_{\rho_{j}}\right)}{\sum_{j=1}^{N}\exp\left(w_{\rho_{j}}\right)},$ (7)
@@
where $\gamma_{\rho_{j}}$ denotes the weight of meta-path $\rho_{j}$. According to the attention mechanism [2], when the semantic attention score $\gamma_{\rho_{j}}$ is higher, it indicates the meta-path $\rho_{j}$ gets more importance in downstream tasks. We aggregate feature embeddings from all meta-paths to obtain the final node features under the meta-path view.
$\boldsymbol{h}_{e}^{2}=\sum_{j=1}^{N}\gamma_{\rho_{j}}\cdot\boldsymbol{h}_{e}^{\rho_{j}}$ (8)
The key idea of generative learning lies in feature reconstruction. Therefore, we fit the reconstructed masked node features $\boldsymbol{h}_{e}^{2}$ and initial features $\boldsymbol{h}_{e}^{0}$, using the Mean Squared Error (MSE) [43] as the loss function:
$\mathcal{L}_{Gen}=\frac{1}{|\mathcal{V}^{[Mask]}|}\sum_{e\in\mathcal{V}^{[Mask]}}\left(\boldsymbol{h}_{e}^{0}-\boldsymbol{h}_{e}^{2}\right)^{2}$ (9)
where $e$ denotes the node in the masked set. Previous GAEs [44] typically used a single-layer MLP or Laplacian matrix in decoding and focused more on restoring the graph structure.
@@
Figure 4: Our GC-HGNN proposed a hard sampling strategy for inter-contrast. Location-aware Pos: the green node uses metapath2vec [3] to perceive the location within a heterogeneous graph to obtain node embeddings. Semantic-aware Pos: The blue node is connected to the anchor node through all types of meta-paths, indicating the semantic similarity between them.
III-C Hierarchical Contrastive Learning
The inputs for this section are the embeddings derived from the network schema and meta-path, denoted as $\boldsymbol{h}{e}^{1}$ and $\boldsymbol{h}{e}^{2}$. To mitigate sampling bias and information loss within the contrastive discriminator, we introduce innovative sampling strategies and a hierarchical contrast framework. The model utilizes two distinct sampling strategies: node-level contrast for intra-contrast and edge-level contrast for inter-contrast. The process of edge-level sampling is depicted in Figure 4. Subsequently, we obtain an edge sampling set, denoted as $\mathcal{E}$.
To obtain high-quality positive samples, we propose two sampling strategies: Location-aware Pos and Semantic-aware Pos, which automatically extract positive samples from the data. As shown in Figure 4, the green nodes represent Location-aware Pos, while the blue nodes denote Semantic-aware Pos.
  • Location-aware Pos calculates the similarity between “$P_{1}-P_{3}$” using node embeddings pre-trained by Metapath2vec [3] and selects the top-k nodes to complete the sampling process. Metapath2vec employs meta-path-based random walks to effectively capture structural information in HGs.
  • Semantic-aware Pos connects “$P_{1}-P_{2}$” through two meta-paths, “$P_{1}-S_{1}-P_{2}$” (PSP) and “$P_{1}-A_{1}-P_{2}$” (PAP), rather than a single meta-path. The more meta-path connections exist, the higher the semantic similarity between them.
We chose metapath2vec for position-aware positives as it captures structural information through meta-path-based walks, making it better suited for long-sequence link awareness. Relying solely on model-based embeddings would weaken position awareness and lose distant node relations [23].
@@ III-C Hierarchical Contrastive Learning
Intra-contrast. First, we utilize sampling set $\mathcal{E}$ to select the positive sample set $\widetilde{\mathcal{V}}$ for node-level contrast based on the degree of nodes. High-degree nodes have more connections in the graph and are more likely to capture the central structure and important patterns in the graph. The embeddings of the model, based on multi-views, are denoted by $\left\{\boldsymbol{H}_{Re}^{\rho_{1}},\ldots,\boldsymbol{H}_{Re}^{\rho_{j}},\ldots,\boldsymbol{H}_{Re}^{\rho_{M}}\right\}$. Next, we take these embeddings as input to obtain a multi-view loss. Intra-contrast employs Scaled Cosine Error (SCE) [10] as the contrastive loss between multi-views. SCE has been proven to focus on hard samples in the data to improve the stability of representation learning.
@@
$\mathcal{L}_{Intra}=\frac{1}{|\widetilde{\mathcal{V}}|}\sum_{v_{i}\in\widetilde{\mathcal{V}}}\left(1-\frac{\boldsymbol{x}_{i}^{T}\boldsymbol{z}_{i}}{\left\|\boldsymbol{x}_{i}\right\|\cdot\left\|\boldsymbol{z}_{i}\right\|}\right)^{\eta},\eta\geq 1$ (10)
@@
Inter-contrast. Inter-contrast performs collaborative contrastive learning between two non-aligned views. In set $\mathcal{E}$, we define the target node to be sampled as $i$ and the connected nodes to it as a positive set $\mathbb{P}_{i}$, inspired by [7]. Otherwise, they form the negative sample set $\mathbb{N}_{i}$. From the scales of $\mathcal{E}$ and $\widetilde{\mathcal{V}}$, inter-contrast typically has more hard negative samples, and thus, its supervisory role is more evident. The contrastive loss of the network schema view with respect to the meta-path view is as follows:
$\mathcal{L}_{Inter}^{1}=-\log\frac{\sum_{j\in\mathbb{P}_{i}}\exp\left(\operatorname{sim}\left(h_{i}^{1},h_{j}^{2}\right)/\tau\right)}{\sum_{k\in\left\{\mathbb{P}_{i}\cup\mathbb{N}_{i}\right\}}\exp\left(\operatorname{sim}\left(h_{i}^{1},h_{k}^{2}\right)/\tau\right)}$ (11)
where $\operatorname{sim}(i,j)$ is the cosine similarity, and $\tau$ denotes a temperature parameter.
Inter-contrast utilizes mutual contrast between the two views to achieve the collaborative process. The contrastive loss of the meta-path view with respect to the network schema view is $\mathcal{L}_{Inter}^{2}$. The overall inter-contrast losses are as follows:
$\mathcal{L}_{{Inter}}=\frac{1}{|\mathcal{E}|}\sum_{i\in\mathcal{E}}\left[\lambda\cdot\mathcal{L}_{Inter}^{1}+(1-\lambda)\cdot\mathcal{L}_{Inter}^{2}\right]$ (12)
where $\lambda$ is a balance coefficient to control the weight of each view pair. We employ intra-contrast and inter-contrast for hierarchical contrastive learning and analyze their operational mechanisms. The intra-contrast assumes that meta-paths are interconnected and aims to maximize their mutual information [14]. This method integrates the different semantics across meta-paths to form a generalized knowledge representation for downstream tasks. On the other hand, the inter-contrast approach attempts to enhance node representation by contrasting local and global information, preventing nodes from being influenced by noise from higher-order neighbors. The hierarchical contrast and sampling strategies are designed to construct a contrastive discriminator to enhance the adversarial interplay between the generative and contrastive models.
III-D Model Optimization